A proton, an electron, and an alpha particle (of charge +2e) form an equilateral triangle of side length 2.51x1010 m. What is the electric potential energy of the three-particle system?-9.179E-8J Submit Answer Incorrect. Tries 3/5 Previous Tries
Answer #1

U = k ( q1 q2 + q2 q3 + q1 q3) / r

U = 9*10^9*10^-38* ( - 1.6*1.6 - 1. 6*2*1.6 + 2* 1.6*1.6) / (2.51*10^-10)

U = - 9.179*10^-19 J

========

do comment in case any doubt, will reply for sure.. Goodluck